OpenStudy (anonymous):

What are the exact solutions of x2 = 4 − 7x? x = x equals negative 7 plus or minus the square root of thirty-three all over 2 x = x equals negative 7 plus or minus the square root of sixty-five all over 2 x = x equals 7 plus or minus the square root of sixty-five all over 2 x = x equals 7 plus or minus the square root of thirty-three all over 2

OpenStudy (anonymous):

\[x^{2}+7x-4=0, x=-7\pm \sqrt{49-4*1*(-4)}, x=-7\pm \sqrt{49+16}, x=-7\pm \sqrt{65}\] so second one

OpenStudy (anonymous):

everything divided by 2a=2*1=2